OpenStudy (anonymous):

a boy throw a ball with 15 m/sec speed at angle 53 and then find distance travel by ball and max height will ball get and time of flight

OpenStudy (anonymous):

@Cubi-Cal @katiebum479 @KissesFromKristine @veeveeniko @Elsa213 @RileyClaim

OpenStudy (anonymous):

@kmullis6

OpenStudy (anonymous):

@katiebum479 do you have ans of this question

OpenStudy (anonymous):

anybody of you have ans please tell me

OpenStudy (anonymous):

I dont even know how to do math like at all im sorry

OpenStudy (anonymous):

I dont know physics either sorry

OpenStudy (anonymous):

@Elsa213 and @Cubi-Cal you both have ans

OpenStudy (anonymous):

Math is not my strongest subject, sorry.

OpenStudy (anonymous):

any of you know ans

OpenStudy (anonymous):

@Somy

OpenStudy (anonymous):

@RileyClaim and @Elsa213

OpenStudy (anonymous):

please tell me any best way to get out of sadness and boreness

OpenStudy (anonymous):

@katiebum479 do you have

OpenStudy (anonymous):

@kisstherains

OpenStudy (anonymous):

I'm not very good with math. :(

OpenStudy (somy):



OpenStudy (anonymous):

@_Army_Strong_

OpenStudy (anonymous):

it is from progetion of object

OpenStudy (quietus):

This is physics btw?? :O

OpenStudy (somy):

hmm there is no time even given hhh

OpenStudy (anonymous):

time is not given you have to find out it

OpenStudy (somy):

i know

OpenStudy (anonymous):

@KissesFromKristine do you have ans

OpenStudy (anonymous):

thanks for try i will be your fan for this try

OpenStudy (anonymous):

Okay I think I have what you need

OpenStudy (anonymous):

if any body want to be my friend you can also be

OpenStudy (somy):

oh i think we can make simultaneous equations

OpenStudy (anonymous):

R = (u^2*sin2x)/g where R is the total distance traveled by the ball thrown, u is the initial velocity and x is the angle to the horizontal the ball was thrown

OpenStudy (anonymous):

g is of course the acceleration of free fall (9.81 m s^-2)

OpenStudy (somy):

i think you can use motion formula \[v^2 = u^2 + 2as\] u= o so \[v^2 = 2as\] u have v= 15m/s and a= 9.81 m/s2 so find S then now u have S - it'll by hypotenuse in the triangle and we need to find opposite we have angle so now find X which is going to be maximum height next use formula \[s= ut + \frac{ 1 }{ 2 }at^2\] again u is 0 so \[s= \frac{ 1 }{ 2 }at^2\] u have S and u know a so find t